OpenStudy (anonymous):

If sin θ = 1/4 and tan θ > 0, what is the value of cos θ?

OpenStudy (anonymous):

A) negative square root of 15 B) square root of 15 C) square root of 15 over 4 D) negative square root of 15 over 4

OpenStudy (nayef):

I will give you a hint: \[\theta is \in the first quadrant. That means \between 0 and 90 degree.\]

OpenStudy (anonymous):

and im lost

OpenStudy (nayef):

hhhha

OpenStudy (nayef):

Okey you know: \[\cos^2\theta+\sin^2\theta=1\]

OpenStudy (anonymous):

sure

OpenStudy (nayef):

\[\cos^2\theta+(\frac{ 1 }{ 4 })^2=1\]

OpenStudy (nayef):

Do you know how to solve the previus eauation!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

look i appreciate ur helping me but i just want the answer plz

OpenStudy (nayef):

You have to learn :)

OpenStudy (nayef):

\[\cos \theta=\sqrt{1-\frac{ 1 }{ 16 }}\]

OpenStudy (nayef):

insert it in the calc!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.96?

OpenStudy (nayef):

\[\frac{ \sqrt{15} }{ 4 }\]

OpenStudy (anonymous):

i should learn to read it says the SINE

OpenStudy (anonymous):



OpenStudy (anonymous):

adjacent sides is \(\sqrt{4^2-1^2}=\sqrt{15}\)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so the cosine is \(\frac{\sqrt{15}}{4}\) same answer as the other way, just a different perspective